# Narrow line AGN selection in CEERS: spectroscopic selection, physical properties, X-ray and radio analysis

### v1

@@ 2.3 Spectral Stacking
Once the final sample of 52 NLAGN has been selected, to get the average emission line properties of the Star-forming galaxies (SFG) and NLAGN samples, we stacked the MR spectra for each of the emission line complexes investigated in this work. The spectral stacking was performed firstly, normalizing all the NLAGN and SFG spectra by the peak flux of the [O iii] 5007 line, in order not to be biased towards the brightest objects. The [O iii] 5007 line is the most common and well-detected line among the CEERS MR spectra, but is not available in 8 AGN and in 14 SFG spectra because for these sources the line falls into a detector gap. Therefore we excluded them from the stacking. We shifted all the spectra to the rest frame, and then we resampled each of the spectra to a uniform and common wavelength grid, given that the three components of each spectrum (corresponding to the three different gratings) have different resolutions and wavelength bins.
The new, and uniform, wavelength grid was defined, for each line complex, by the wavelength bin allowing for the best spectral resolution among the different resolutions of the spectra involved in the stacking of the line(s). We checked that the results do not significantly change taking the wavelength bin corresponding to the worst spectral resolution among the spectra. Before stacking the spectra, we also fit and subtract the continuum from the rest-frame rebinned spectra.
@@
The resulting spectra from the stacking of all NLAGN and SFG are presented in Fig. 1.
@@ 2.4 SED fitting
To investigate the physical properties of the host galaxies of the selected NLAGN and to compare them with the properties of the parent sample, we performed a SED fitting analysis using CIGALE (Boquien et al., 2019). Starting from the 217 spectra of the MR sample with a secure redshift, we first cross-match the target coordinates with the HST+JWST publicly available photometric catalog of the CEERS field reported in the DJA archive222https://s3.amazonaws.com/grizli-v2/JwstMosaics/v7/index.html. Considering a 0.5” matching radius we found 117 counterparts. This catalog includes seven HST bands (F435W, F606W, F814W, F105W, F125W, F140W, F160W) and ten JWST/NIRCam and MIRI bands (F115W, F150W, 182M, F200W, 210M, F277W, F356W, F410M, F444W, F770W). Since not the whole CEERS MR sample is covered by JWST imaging, we also cross-match the remaining sources with the 3D-HST multiwavelength catalog (Momcheva et al., 2016; Skelton et al., 2014), covering the EGS field and using the same cross-matching radius to avoid false matches. We found 93 additional counterparts, bringing the total number of sources with an associated optical/NIR photometry to 210. For the remaining 7 spectra, we didn’t find a counterpart closer than 0.5”, but their counterparts are returned when we considered a larger cross-matching radius (). For sources detected in the JWST+HST catalog, we didn’t consider the additional photometry of the 3D-HST catalog since most of these sources are undetected in photometric bands bluer than the HST bands (being at high-), while the photometry at longer wavelengths than HST bands is already covered by JWST (the deepest in the near and mid-infrared).
@@
We used delayed star formation history (SFH) models for both redshift groups because they are able to reproduce both early-type and late-type galaxies, with an additional term that allows for a recent (and constant) variation of the star-formation rate (SFR) that can be in the burst or in the quench phase. We adopted stellar templates from Bruzual & Charlot (2003), and a Chabrier initial mass function (Chabrier, 2003). We also include the nebular emission module that is extremely important to account for the contribution of emission lines in the broad-band photometry (Schaerer & de Barros, 2012; Salvato et al., 2019), in particular for the low-mass and high- regimes, where emission lines can account for a consistent fraction of the wideband photometric flux. This module is computed by CIGALE based on a grid of user-provided parameters and in a self-consistent manner using a grid of Cloudy photoionization calculations Ferland et al. (2013). For the attenuation of the stellar continuum emission, we considered the dustatt_modified_CF00 module (Charlot & Fall, 2000), which allows different attenuations for the young and old stellar populations. We also include dust emission in the IR following the empirical templates of Dale et al. (2014). The main differences in the grid with respect to the low- one consist in the larger parameter space explored for the possible final starburst state of high- galaxies and in the lower metallicities and higher ionization parameters allowed in the stellar and nebular modules, according with recent JWST results (Endsley et al., 2023; Tang et al., 2023). Indeed, SFG at high- are observed to be more frequent in a bursty SF regime (Faisst et al., 2020; Dressler et al., 2023; Looser et al., 2023) and also, the gas metallicities are generally 0.5-1 dex lower than in the local universe (Curti et al., 2023b; Nakajima et al., 2023).
@@
For sources classified as AGN based on our NLAGN selection, as well as the BLAGN selected in Harikane et al. (2023), we included AGN modules in the SED fitting to account for the AGN emission. We employed the skirtor2016 module introduced in Yang et al. (2020), which is widely used in the community and has demonstrated reliability in studying various aspects of AGN (e.g., Mountrichas et al., 2022; López et al., 2023; Yang et al., 2023b). The SED produced by the AGN combines emissions from the accretion disk, torus, and polar dust. The accretion disk, responsible for the UV-optical emission at the central region, is parameterized according to Schartmann et al. (2005), with a typical delta value of -0.36. Photons from the accretion disk can be obscured and scattered by dust in the vicinity, within the torus and/or in the polar direction. Specifically, for the torus, the skirtor2016 module employs a clumpy two-phase model (Stalevski et al., 2016), based on the 3D radiative-transfer code SKIRT (Baes et al., 2011; Camps & Baes, 2015). As our photometry does not cover the restframe mid-IR, we fixed the optical depth at 9.7 m at = 3. For the polar dust component, we used the Small Magellanic Cloud (SMC) extinction curve, recommended due to its preference in AGN observations (e.g., Bongiorno et al., 2012). The extinction amplitude, parameterized as , is a user-defined free parameter for which we chose a typical value of . Emission from the polar dust maintains energy conservation, assuming isotropic emission and a ‘grey body’ model (Casey, 2012). Finally, we allowed inclination values from 50 to 80 degrees and varied the AGN fraction (defined as the ratio between the AGN luminosity and the host galaxy luminosity between 0.15 and 2 microns) from 0.1 (weak AGN contamination) to 0.7 (dominant AGN emission).
@@ 2.5 X-ray and radio stacking
To investigate the multiwavelength properties of the CEERS NLAGN selected using the diagnostic diagrams, we look at their counterparts in the X-rays and Radio images of the EGS field.
@@
The Chandra AEGIS-XD field is the third deepest X-ray field ever performed, with an exposure that reaches 800 ks. The X-ray observations and the related X-ray catalogue are described in Nandra et al. (2015). The X-ray catalogue contains 937 X-ray sources, and 553 of these sources (those with enough photon counts) also have an X-ray spectral analysis performed by Buchner et al. (2015). By cross-matching the CEERS MR catalog with the X-ray catalogue, we found 7 matches, involving sources at . The X-ray spectral analysis classified two of these sources as galaxies (CEERS-3060, CEERS-3051), the other 5 (CEERS-2919, CEERS-2808, CEERS-2904, CEERS-2989, CEERS-2900) as AGN. In particular, CEERS-2904, CEERS-2989, and CEERS-2919 are part of the new low-redshift BLAGN sample presented in Appendix A.
@@
Four out of the five X-ray AGN were selected as NLAGN by our spectroscopic selection, as we are going to present in Sec. 3. However, all the other 48 selected NLAGN do not show any indication of X-ray emission. Therefore, we decided to perform an X-ray stacking analysis of the NLAGN not X-ray detected using the publicly available code CSTACK v4.5 Miyaji et al. (2008). For each target, CSTACK provides the net (background-subtracted) count rate in the soft (0.5–2 keV) and hard (2–8 keV) bands using all the observations from the AEGIS-XD survey and the associated exposure maps. Since the survey is a mosaic, multiple observations at different off-axis angles can cover the same position. Due to the variation of the Chandra PSF with the off-axis angle, for each observation of an object, CSTACK defines a circular source extraction region with the size determined by the 90% encircled counts fraction (ECF) radius () (with a minimum of 1”), to optimize the signal-to-noise ratio of the stacked signals. We set the background region for each source to be a arcsec2 area centered on the object, excluding the region around the object used to compute the net counts. Moreover, the code excludes the photon counts of the background circular regions around all the detected X-ray sources, with radii that depend on the net counts of the X-ray source.
@@
We derived the count rates in the soft band (0.5-2keV, SB), the band for which the detector is most sensitive and can, therefore, provide tighter constraints on the X-ray analysis. Then we converted the count rates into fluxes assuming an X-ray power spectrum with an intrinsic spectral index and using the Chandra tool PIMMS333https://cxc.harvard.edu/toolkit/pimms.jsp.
@@
The field targeted by the CEERS survey is also covered by 1.4GHz observations, the AEGIS20 described in Ivison et al. (2007). This radio survey was performed in 2006 with the Very Large Array (VLA) over an area of deg2 and reaching, in the central region, a rms of Jy. The resulting radio catalog (Ivison et al., 2007) contains 1123 individual radio sources, of which only two are part of the CEERS MR sample, i.e., CEERS-2900 and CEERS-3129, the first (that is also X-ray detected) showing the typical morphology of a radio-loud AGN. Both these sources were identified as AGN also by our selection, but all the other NLAGN do not have any radio counterparts. Therefore, we perform radio stacking analysis to investigate the radio statistical properties of the NLAGN population below the survey detection threshold. We applied a pixel-by-pixel median stacking, taking arcsec2 cutouts around the sources we wanted to stack. This method was proved (White et al., 2007) to be more robust (compared to the mean stacking, for example) towards systematic effects (i.e. confusion) and also less affected by possible high-flux density contaminants. Since the CEERS survey covers only a very small region of the radio image on which the rms is almost uniform, we do not weigh each radio cutout for the corresponding noise value in the stacking procedure. We also estimated the final rms of the stacking images using a median absolute deviation procedure, following Keller et al. (2024).
@@ 3 Results
Here we present the result of the NLAGN selection performed on the sample of the 217 MR CEERS spectra with a secure redshift identification. The NLAGN selection was performed using emission line diagnostic diagrams.
In all the diagrams, we plot both the observed data points and also the line ratios derived from the photoionization models described in Feltre et al. (2016) and Gutkin et al. (2016). These models were computed using the Cloudy code (Ferland et al., 2013) for SFG and AGN NL regions and for various gas metallicities, ionization parameters, dust content, and ISM densities, and considering a wide range of the parameters space (for the full grid of values covered by the parameter space we refer to Table 1 in Feltre et al., 2016). To allow a comprehensive enough coverage of the SFG models but to avoid physical conditions rarely (or never) found in the general populations of SFG in the local or high- Universe, we applied to these models the same cut in the parameter space as presented and discussed in Scholtz et al. (2023a).
@@
We also mark with distinctive colours the seven sources that are X-ray detected (gold), the two radio-detected sources (green), and the eight sources that were selected as BLAGN at by Harikane et al. (2023) (red). In each diagnostic diagram, the sources with magenta edges are those selected as NLAGN in that specific diagnostic.
It is worth noting that the line fluxes involved in the diagnostic diagrams described in the next Sections are not corrected for the effect of dust. This is because we are always considering ratios between lines close to each other, hence subjected to almost the same reddening, making the ratio insensible to the presence of dust. The few exceptions are discussed and justified. The acronyms used for the line ratios are reported in Tab. 1.
@@ 3.1 Optical diagnostics
In Fig. 2, we show the traditional BPT (Baldwin et al., 1981), the S2-VO87, and the R3O1 (Veilleux & Osterbrock, 1987) diagnostics diagrams based on the R3-N2, R3-S2, and R3-O1 line ratios, respectively.
@@
The traditional demarcation lines between SFG and AGN provided by Kauffmann et al. (2003) and Kewley et al. (2001) already proved to be way less effective in selecting high- AGN compared to the local universe (Scholtz et al., 2023a; Übler et al., 2023; Maiolino et al., 2023). Indeed, the lower metallicities of high- sources make [N ii] (but also [S ii] ) emission lines fainter, shifting the objects (AGN included) towards the left part of the BPT diagram. At the same time, the higher ionization parameter of the high- SFG (Cameron et al., 2023; Sanders et al., 2023; Topping et al., 2024), due to, on average, the presence of younger stars and lower metallicities compared to the local universe, moves the SFG towards higher R3 ratios, generating a large overlap between AGN and SFGs. Therefore, in the R3N2 and R3S2 diagnostics, we used the conservative demarcation lines provided in Scholtz et al. (2023a), derived considering the distribution of the photoionization models of Gutkin et al. (2016); Feltre et al. (2016) and Nakajima & Maiolino (2022).
@@
Since the R3O1 diagnostic diagram was not included in the analysis of Scholtz et al. (2023a), we followed the same fitting method to the photoionization models to find more conservative demarcation lines with respect to those presented in Kewley et al. (2001). We derived the following dividing line between the AGN and SFG population:
@@
As it is possible to see from the three plots, none of the BLAGN at high- are in the AGN region of the three diagnostics, based on the new demarcation lines, and only one would be selected considering the traditional demarcation lines in the R3S2 and R3O1 diagrams (is the same source, CEERS-1665).
@@
The five AGNs selected with the R3N2 diagram are all at . Three of them are X-ray-detected AGN (one is also radio-detected, i.e. CEERS-2900), and one is another radio-detected source (CEERS-3129). It is worth noting that even the X-ray source at the highest redshift (i.e. CEERS-2808 at ), does not fall in the pure-AGN region of this diagnostic diagram.
@@
The R3S2 diagnostic diagram appears effective in selecting AGN, also at a higher redshift with respect to the BPT. The new demarcation line provided in Scholtz et al. (2023a) partially overlaps with the traditional one proposed in Kewley et al. (2001). In this case, we select 21 AGN, six of which at , one at . None of the BLAGN is selected as AGN based on this diagnostic, while three X-ray sources and the two radio sources are confirmed as AGN also in this diagnostic.
@@
Also the R3O1 diagnostic diagram is more effective in identifying NLAGN compared to the R3N2. In this diagram, we select 12 AGN, 7 at (including 2 X-ray sources), and 5 at , one of these at . We will further discuss the effectiveness of this diagnostic and of the R3S2 in Sec. 4.1.
@@
In Fig. 3 we show the diagnostic based on the He2 vs N2 lines ratio. In this case, the demarcation line is the original one provided by Shirazi & Brinchmann (2012), which still holds even at high-, as also reported in Dors et al. (2024) and Tozzi et al. (2023). Overall, we detect He ii 4686 in seven sources: all except one of these are classified as AGN. None of the CEERS BLAGN show detection of the He ii 4686 line. In this diagnostic, there are also five sources that are selected as AGN only based on the N2 ratio.
@@
The last diagnostics exploiting rest-frame optical lines are the two diagnostics based on the O3Hg ratio and proposed in Mazzolari et al. (2024). Hereafter, we will refer with M1 to the O3Hg versus O32 diagnostic diagram and with M2 to the O3Hg versus Ne3O2 diagram. These diagnostics proved to be effective in selecting those AGN characterized by high O3Hg ratios at a given ratio of O32 or Ne3O2, and were able to select NLAGN also at (Mazzolari et al., 2024).

### v3

@@ 2.3 Spectral stacking
Once the final sample of 52 NLAGNs has been selected, to get the average emission line properties of the SFGs and NLAGN samples, we stacked the MR spectra for each of the emission line complexes investigated in this work. The spectral stacking was performed by normalizing all the NLAGN and SFG spectra by the peak flux of the [O iii] 5007 line in order not to be biased toward the brightest objects. The [O iii] 5007 line is the most common and well-detected line among the CEERS MR spectra, but is not available in 8 AGNs and in 14 SFG spectra because for these sources the line falls into a detector gap. Therefore, we excluded them from the stacking. We shifted all the spectra to the rest frame, and then we resampled each of the spectra to a uniform and common wavelength grid, given that the three components of each spectrum (corresponding to the three different gratings) have different resolutions and wavelength bins.
The new and uniform wavelength grid was defined, for each line complex, by the wavelength bin allowing for the best spectral resolution among the different resolutions of the spectra involved in the stacking of the line(s). We checked that the results do not significantly change, taking the wavelength bin corresponding to the worst spectral resolution among the spectra. Before stacking the spectra, we also fit and subtracted the continuum from the rest-frame rebinned spectra.
@@
The resulting spectra from the stacking of all NLAGNs and SFGs are presented in Fig. 1.
@@ 2.4 SED fitting
To investigate the physical properties of the host galaxies of the selected NLAGNs and to compare them with the properties of the parent sample, we performed a SED fitting analysis using CIGALE (boquien19). Starting from the 217 spectra of the MR sample with a secure redshift, we first crossmatch the target coordinates with the HST+JWST publicly available photometric catalog of the CEERS field reported in the DJA archive333https://s3.amazonaws.com/grizli-v2/JwstMosaics/v7/index.html. Considering a 0.5” matching radius we found 117 counterparts. This catalog includes seven HST bands (F435W, F606W, F814W, F105W, F125W, F140W, F160W) and ten JWST/NIRCam and MIRI bands (F115W, F150W, 182M, F200W, 210M, F277W, F356W, F410M, F444W, F770W). Since not the whole CEERS MR sample is covered by JWST imaging, we also crossmatched the remaining sources with the 3D-HST multiwavelength catalog (Momcheva16; Skelton14), covering the EGS field and using the same crossmatching radius to avoid false matches. We found 93 additional counterparts, bringing the total number of sources with an associated optical/NIR photometry to 210. For the remaining seven spectra, we did not find a counterpart closer than 0.5”, but their counterparts are returned when we considered a larger crossmatching radius (). For sources detected in the JWST+HST catalog, we did not consider the additional photometry of the 3D-HST catalog since most of these sources are undetected in photometric bands bluer than the HST bands (being at high-), while the photometry at longer wavelengths than HST bands is already covered by JWST (the deepest in the near and mid-infrared).
@@
We used delayed star formation history (SFH) models for both redshift groups because they are able to reproduce both early-type and late-type galaxies, with an additional term that allows for a recent (and constant) variation in the star formation rate (SFR) that can be in the burst or in the quench phase. We adopted stellar templates from bruzual03, and a Chabrier initial mass function (chabrier03). We also include the nebular emission module that is extremely important to account for the contribution of emission lines in the broad-band photometry (Schaerer2012; Salvato2019), in particular for the low-mass and high- regimes, where emission lines can account for a consistent fraction of the wideband photometric flux. This module is computed by CIGALE based on a grid of user-provided parameters and in a self-consistent manner using a grid of Cloudy photoionization calculations Ferland13. For the attenuation of the stellar continuum emission, we considered the dustatt_modified_CF00 module (Charlot00), which allows different attenuations for the young and old stellar populations. We also include dust emission in the IR following the empirical templates of Dale14. The main differences in the grid with respect to the low- one consist in the larger parameter space explored for the possible final starburst state of high- galaxies and in the lower metallicities and higher ionization parameters allowed in the stellar and nebular modules, according with recent JWST results (Endsley23; Tang23). Indeed, SFGs at high- are observed to be more frequent in a bursty SF regime (Faisst20; Dressler23; Looser23) and also, the gas metallicities are generally 0.5-1 dex lower than in the local Universe (Curti23b; Nakajima23).
@@
For sources classified as AGNs based on our NLAGN selection, as well as the BLAGNs selected in Harikane23, we included AGN modules in the SED fitting to account for the AGN emission. We employed the skirtor2016 module introduced in yang20, which is widely used in the community and has demonstrated reliability in studying various aspects of AGN (e.g., Mountrichas22; LopezIE23; Yang2023). The SED produced by the AGN combines emissions from the accretion disk, torus, and polar dust. The accretion disk, responsible for the UV-optical emission at the central region, is parameterized according to Schartmann05, with a typical delta value of -0.36. Photons from the accretion disk can be obscured and scattered by dust in the vicinity, within the torus and/or in the polar direction. Specifically, for the torus, the skirtor2016 module employs a clumpy two-phase model (Stalevski16), based on the 3D radiative-transfer code SKIRT (SKIRTBaes2011ApJS..196...22B; SKIRTCamps2015A&C.....9...20C). As our photometry does not cover the restframe mid-IR, we fixed the optical depth at 9.7 m at = 3. For the polar dust component, we used the Small Magellanic Cloud (SMC) extinction curve, recommended due to its preference in AGN observations (e.g., 2012MNRAS.427.3103B). The extinction amplitude, parameterized as , is a user-defined free parameter for which we chose a typical value of . Emission from the polar dust maintains energy conservation, assuming isotropic emission and a “gray body” model (2012MNRAS.425.3094C). Finally, we allowed inclination values from 50 to 80 degrees and varied the AGN fraction (defined as the ratio between the AGN luminosity and the host galaxy luminosity between 0.15 and 2 microns) from 0.1 (weak AGN contamination) to 0.7 (dominant AGN emission).
@@ 2.5 X-ray and radio stacking
To investigate the multiwavelength properties of the CEERS NLAGNs selected using the diagnostic diagrams, we looked at their counterparts in the X-rays and Radio images of the EGS field.
@@
The Chandra AEGIS-XD field is the third deepest X-ray field ever performed, with an exposure that reaches 800 ks. The X-ray observations and the related X-ray catalog are described in nandra15. The X-ray catalog contains 937 X-ray sources, and 553 of these sources (those with enough photon counts) also have an X-ray spectral analysis performed by buchner15. By crossmatching the CEERS MR catalog with the X-ray catalog, we found seven matches, involving sources at . The X-ray spectral analysis classified two of these sources as galaxies (CEERS-3060, CEERS-3051), the other 5 (CEERS-2919, CEERS-2808, CEERS-2904, CEERS-2989, CEERS-2900) as AGN. In particular, CEERS-2904, CEERS-2989, and CEERS-2919 are part of the new low-redshift BLAGN sample presented in Appendix A.
@@
Four out of the five X-ray AGNs were selected as NLAGNs by our spectroscopic selection, as we are going to present in Sec. 3. However, none of the other 48 NLAGNs selected show any indication of X-ray emission. Therefore, we decided to perform an X-ray stacking analysis of the NLAGNs not X-ray detected using the publicly available code CSTACK v4.5 miyaji08. For each target, CSTACK provides the net (background-subtracted) count rate in the soft (0.5–2 keV) and hard (2–8 keV) bands using all the observations from the AEGIS-XD survey and the associated exposure maps. Since the survey is a mosaic, multiple observations at different off-axis angles can cover the same position. Due to the variation in the Chandra PSF with the off-axis angle, for each observation of an object, CSTACK defines a circular source extraction region with the size determined by the 90% encircled counts fraction radius () (with a minimum of 1”), to optimize the S/N of the stacked signals. We set the background region for each source to be a arcsec2 area centered on the object, excluding the region around the object used to compute the net counts. Moreover, the code excludes the photon counts of the background circular regions around all the detected X-ray sources, with radii that depend on the net counts of the X-ray source.
@@
We derived the count rates in the soft band (0.5-2keV, SB), the band for which the detector is most sensitive and can, therefore, provide tighter constraints on the X-ray analysis. Then we converted the count rates into fluxes assuming an X-ray power spectrum with an intrinsic spectral index and using the Chandra tool PIMMS444https://cxc.harvard.edu/toolkit/pimms.jsp.
@@
The field targeted by the CEERS survey is also covered by 1.4GHz observations, the AEGIS20 described in Ivison07. This radio survey was performed in 2006 with the Very Large Array (VLA) over an area of deg2 and reaching, in the central region, a rms of Jy. The resulting radio catalog (Ivison07) contains 1123 individual radio sources, of which only two are part of the CEERS MR sample, CEERS-2900 and CEERS-3129, the first (that is also X-ray detected) showing the typical morphology of a radio-loud AGN. Both these sources were also identified as AGNs by our selection, but none of the other NLAGNs have any radio counterparts. Therefore, we performed a radio stacking analysis to investigate the radio statistical properties of the NLAGN population below the survey detection threshold. We applied a pixel-by-pixel median stacking, taking arcsec2 cutouts around the sources we wanted to stack. This method was proved (white07) to be more robust (compared to the mean stacking, for example) toward systematic effects (i.e., confusion) and also less affected by possible high-flux density contaminants. Since the CEERS survey covers only a very small region of the radio image on which the rms is almost uniform, we did not weigh each radio cutout for the corresponding noise value in the stacking procedure. We also estimated the final rms of the stacking images using a median absolute deviation procedure, following Keller24.
@@ 3 Results
Here, we present the result of the NLAGN selection performed on the sample of the 217 MR CEERS spectra with a secure redshift identification. The NLAGN selection was performed using emission line diagnostic diagrams.
In all the diagrams, we plot both the observed data points and also the line ratios derived from the photoionization models described in Feltre16 and Gutkin16. These models were computed using the Cloudy code (Ferland13) for SFG and AGN NL regions and for various gas metallicities, ionization parameters, dust content, and ISM densities, and considering a wide range of the parameters space (for the full grid of values covered by the parameter space we refer to Table 1 in Feltre16). To allow for a sufficiently comprehensive coverage of the SFG models but to avoid physical conditions rarely (or never) found in the general populations of SFG in the local or high- Universe, we applied to these models the same cut in the parameter space as presented and discussed in Scholtz23b. In particular, we selected all models with metallicities between 0.001 and 0.02 (corresponding to 0.06-1.3 solar) and a dust-to-metal mass ratio of 0.3, which is intermediate between the range observed in the most metal-poor absorbers (Konstantopoulou23) and the Milky-Way value of 0.45. We consider all models with carbon to oxygen abundance ratio in the range 0.38-1.00 solar to describe a variety of different or less common star formation grids.
@@
We also mark with distinctive colors the seven sources that are X-ray detected (gold), the two radio-detected sources (green), and the eight sources that were selected as BLAGNs at by Harikane23 (red). In each diagnostic diagram, the sources with magenta edges are those selected as NLAGN in that specific diagnostic.
It is worth noting that the line fluxes involved in the diagnostic diagrams described in the next Sections are not corrected for the effect of dust. This is because we are always considering ratios between lines close in wavelength to each other, hence subjected to almost the same reddening, making the ratio insensible to the presence of dust. The few exceptions are discussed and justified. The acronyms used for the line ratios are reported in Tab. 1.
@@ 3.1 Optical diagnostics
In Fig. 2, we show the traditional BPT diagram (Baldwin81), the R3S2 (also called VO87), and the R3O1 (Veilleux87) diagnostics diagrams based on the R3-N2, R3-S2, and R3-O1 line ratios, respectively.
@@
The traditional demarcation lines between SFG and AGN provided by Kauffmann03 and Kewley01 already proved to be way less effective in selecting high- AGNs compared to the local Universe (Scholtz23b; ubler23; Maiolino23c). Indeed, the lower metallicities of high- sources make [N ii] (but also [S ii] ) emission lines fainter, shifting the objects (AGN included) toward the left part of the BPT diagram. At the same time, the higher ionization parameter of the high- SFG (Cameron23; Sanders23; topping24), due to, on average, the presence of younger stars and lower metallicities compared to the local Universe, moves the SFG toward higher R3 ratios, generating a large overlap between AGN and SFGs. Therefore, in the R3N2 and R3S2 diagnostics, we used the conservative demarcation lines provided in Scholtz23b, derived considering the distribution of the photoionization models of Gutkin16; Feltre16 and Nakajima22.
@@
Since the R3O1 diagnostic diagram was not included in the analysis of Scholtz23b, we followed the same fitting method to the photoionization models to find more conservative demarcation lines with respect to those presented in Kewley01. We derived the following dividing line between the AGN and SFG population:
@@
As it is possible to see from the three plots, none of the BLAGNs at high- are in the AGN region of the three diagnostics, based on the new demarcation lines, and only one would be selected considering the traditional demarcation lines in the R3S2 and R3O1 diagrams (is the same source, CEERS-1665).
@@
The five AGNs selected with the R3N2 diagram are all at . Three of them are X-ray-detected AGNs (one is also radio-detected, i.e., CEERS-2900), and one is another radio-detected source (CEERS-3129). It is worth noting that even the X-ray source at the highest redshift (i.e., CEERS-2808 at ), does not fall in the pure-AGN region of this diagnostic diagram.
@@
The R3S2 diagnostic diagram appears effective in selecting AGN also at a higher redshift compared to the BPT. The new demarcation line provided in Scholtz23b partially overlaps with the traditional one proposed in Kewley01. In this case, we select 21 AGNs, six of which at , one at . None of the BLAGNs is selected as AGN based on this diagnostic, while three X-ray sources and the two radio sources are confirmed as AGN also in this diagnostic.
@@
Also the R3O1 diagnostic diagram is more effective in identifying NLAGN compared to the R3N2. In this diagram, we select 12 AGNs, 7 at (including 2 X-ray sources), and 5 at , one of these at . We further discuss the effectiveness of this diagnostic and of the R3S2 in Sec. 4.1.
@@
In Fig. 3, we show the diagnostic based on the He2 versus N2 lines ratio. In this case, the demarcation line is the original one provided by Shirazi12, which still holds even at high-, as also reported in Dors23 and Tozzi23. Overall, we detect He ii 4686 in seven sources: all except one of these are classified as AGN. None of the CEERS BLAGNs show detection of the He ii 4686 line. In this diagnostic, there are also five sources that are selected as AGN only based on the N2 ratio.
@@
The last diagnostics exploiting rest-frame optical lines are the two diagnostics based on the O3Hg ratio and proposed in Mazzolari24b. Hereafter, we refer with M1 to the O3Hg versus O32 diagnostic diagram and with M2 to the O3Hg versus Ne3O2 diagram. These diagnostics proved to be effective in selecting those AGN characterized by high O3Hg ratios at a given ratio of O32 or Ne3O2, and were able to select NLAGN also at (Mazzolari24b).
@@
We note that in the diagnostics presented in Fig. 2, 3, 4 there are some sources lying close enough to the demarcation lines that their line ratios’ error can potentially place them outside from the AGN-only region of the diagrams. We considered all the selected NLAGN whose 1 uncertainties are compatible with SF-driven photoionization, and we double-checked if they were safely identified as NLAGN in other diagnostics (i.e., without 1 errors crossing the demarcation line). We identified four NLAGNs with errors compatible with SFG ionization – CEERS-2668, CEERS-3535, CEERS-1836, and CEERS-3585 – and we conservatively decided to mark them only as candidates NLAGNs. CEERS-2900 should also be included in this sample, but given its detection in both the X-ray and radio image of the field it can be safely considered a NLAGN.